# Formal Foundations for Controlled Stochastic Activity Networks

### v1

@@ 6 Conclusion
This paper has introduced Controlled Stochastic Activity Networks (Controlled SANs) as a formal, automata-theoretic framework that unifies control, nondeterminism, probability, and stochastic timing within a single modeling language. By developing a layered semantic hierarchy—from controlled automata to probabilistic and stochastic extensions—we have established a rigorous foundation for representing and analyzing distributed real-time systems operating under uncertainty.
Our results formally connect Controlled SANs with classical models such as CTMDPs, GSMPs, and controlled probabilistic automata, while also introducing a structured taxonomy of control policies and corresponding language hierarchies. These theoretical developments enable precise reasoning about expressiveness, equivalence, and compositionality.
A distinctive aspect of this work is its compatibility with AI-driven control synthesis. The framework naturally supports PAC-based reasoning, reinforcement-learning approaches, and other algorithmic techniques for scalable decision making under uncertainty. This bridges classical formal verification with modern AI methods, creating new opportunities for practical verification and adaptive control in complex systems.
Future work will focus on integrating temporal-logic model checking, controller synthesis, and PAC-optimization methods into automated tools, as well as exploring applications in safety-critical domains such as autonomous systems, dependable AI, and cyber-physical infrastructures. By combining rigorous formal semantics with learning-based control, Controlled SANs provide a mathematically principled foundation for the next generation of verifiable intelligent systems.
@@ Appendix A Proofs
Consider the class of activity networks [movaghar2001stochastic] that have only instantaneous activities, standard gates, and a special type of gates called “inhibitor” gates. An inhibitor gate is an input gate with an enabling predicate $g$ and an identity function such that $g(x)=true$ iff $x=0$. The above class of activity networks corresponds to the class of extended Petri nets [Peterson1981]. Using the notion of the register machines of Shepardson and Sturgis [Shepardson1963] or the program machines of Minsky [Minsky1967], it can easily be shown that an extended Petri net can simulate a Turing machine [Hack1975, Peterson1981]. Following a similar approach, one can show that an extended Petri net can, indeed, simulate a nondeterministic Turing machine [Sipser2013] and, hence, can represent any computable relation.
@@
It follows that an activity network with some instantaneous activities, standard gates, and inhibitor gates is, likewise, able to simulate a nondeterministic Turing machine and, hence, can represent any computable relation. This latter result will be used below to prove the theorem.
@@
Let $S=(Q,A,C,\rightarrow,Q_{0})$ be a computable controlled automaton. Without loss of generality, let $Q={\mathcal{N}}$. Define:
  • for $a\in A$ and $c\in C$, $R_{ac}=\{(i,j);i,j\in{\mathcal{N}},i\stackrel{{\scriptstyle a,c}}{{\rightarrow}}j\}$
  • for $a\in A$, $G_{a}:{\mathcal{N}}\longrightarrow\{true,false\}$, where $G_{a}(i)=true$ iff there exists $c\in C$ and $j\in{\mathcal{N}}$ such that $i\stackrel{{\scriptstyle a,c}}{{\rightarrow}}j$,
  • $R_{Q_{0}}=\{(1,i);i\in Q_{0}\}$.
Since $S$ is a computable controlled automaton, $G_{a},R_{ac}$, for $a\in A$ and $c\in C$, and $R_{Q_{0}}$ above will also be computable relations, which means that some nondeterministic Turing machines can represent them. Accordingly, using the result mentioned earlier, $G_{a},R_{ac}$, for $a\in A$ and $c\in C$, and $R_{Q_{0}}$ can be simulated by some activity networks $K_{G_{a}},K_{R_{ac}}$, for $a\in A$ and $c\in C$, and $K_{Q_{0}}$, respectively, which have only some instantaneous activities, standard gates and inhibitor gates.
Now, consider a controlled activity network $K$ with the set of timed activities $A$ and the set of contol actions $C$ such that, for any $a\in A$ and $c\in C$, $K$ includes an controlled activity subnetwork as depicted in Figure 5. In Figure 2, $K_{G_{a}}$ and $K_{R_{ac}}$ represent some controlled activity networks which simulate $G_{a}$ and $R_{ac}$, respectively. $P_{1a}$, $P_{2a}$, $P_{3ac}$, and $P_{M}$ are places such that when $P_{1a}$ is empty all instantaneous activities of $K_{G_{a}}$ are disabled and when $P_{3ac}$ is empty all instatntaneous activities of $K_{R_{ac}}$ are disabled. Initially, $P_{1a}$ has a token but $P_{2a}$ and $P_{3ac}$ are empty. As soon as $P_{S}$ acquires a marking $x$ such that $G_{a}(x)=true$, $K_{G_{a}}$ starts execution and after a finite number of instantaneous activity completions $P_{1a}$ loses a token, $P_{2a}$ gains one, and the marking of $P_{S}$ remains the same. When $P_{3ac}$ obtains a token, $K_{R_{ac}}$ also begins execution and after a finite number of instantaneous activity completions $P_{3a}$ loses a token, $P_{1a}$ gains one, and the marking of $P_{S}$ changes from $x$ to $y$ where $(x,y)\in R_{ac}$. $K$ also includes a controlled activity subnetwork as depicted in Figure 6. In Figure 3, $K_{Q_{0}}$ represents a controlled activity network which simulates $R_{Q_{0}}$. $P_{0}$ is a place with an initial marking of 1. Initially, $P_{S}$ is empty and $K_{Q_{0}}$ starts execution. After a finite number of instantaneous activity completions $P_{0}$ loses a token, the marking of $P_{S}$ is set to $x$ where $(1,x)\in R_{Q_{0}}$, and all activities of $K_{Q_{0}}$ are disabled. Suppose $K$ includes no controlled activity subnetwork other than those depicted in Figure 5 and Figure 6. It can be shown that the controlled activity network $K$ defined above, indeed, realizes a controlled automaton which is isomorphic to $S$ mentioned earlier, and the proof is complete. ∎

### v2

@@ 6 Conclusion
This paper introduced Controlled Stochastic Activity Networks (Controlled SANs) as a unified, automata-theoretic framework that integrates nondeterministic control, probabilistic branching, and stochastic timing into a cohesive modeling language for real-time systems.
By establishing formal connections between Controlled SANs, classical models like CTMDPs, GSMPs, and GSMDPs, and recent extensions such as Controlled Probabilistic Automata, we provide a rigorous semantic hierarchy for analyzing expressiveness and optimality. We also introduced a structured taxonomy of control policies and developed uniform techniques for approximating $\varepsilon$-optimal policies via discretized GSMDPs.
A distinctive contribution of our framework is its compatibility with AI-based control synthesis, including PAC learning, reinforcement learning, and approximate dynamic programming. This creates a robust bridge between traditional formal methods and scalable AI planning under uncertainty.
Future work will focus on incorporating temporal-logic specifications, controller synthesis under logical constraints, and model-checking algorithms into CSA-based tools. Applications in cyber-physical systems, autonomous robotics, and dependable AI are particularly promising. Ultimately, Controlled SANs offer a principled foundation for building the next generation of adaptive, formally verifiable intelligent systems.
@@ Appendix A Proofs
We begin by considering a restricted class of activity networks [49] consisting only of instantaneous activities, standard gates, and a special class of gates known as inhibitor gates. An inhibitor gate is an input gate with an enabling predicate $g$ and the identity function as its output function, where $g(x)=\mathit{true}$ if and only if $x=0$.
This class of activity networks is equivalent, at the structural level, to the class of extended Petri nets with inhibitor arcs [54]. Such nets strictly generalize ordinary Petri nets and form the basis of several expressive stochastic Petri net models, including Generalized Stochastic Petri Nets (GSPNs) [40, 4] when timing is added.
It is a classical result that extended Petri nets with inhibitor arcs can simulate Turing-complete models of computation. Using register machines [64] or Minsky’s program machines [46], one can show that extended Petri nets simulate deterministic and nondeterministic Turing machines [23, 54, 65]. Consequently, they can represent any computable relation.
@@
Since activity networks with instantaneous activities, standard gates, and inhibitor gates are behaviorally equivalent to extended Petri nets, it follows that such activity networks can also simulate a nondeterministic Turing machine and hence represent any computable relation. This universality property forms the basis of the proof.
Let
$S=(Q,A,C,\rightarrow,Q_{0})$
@@
be a computable controlled automaton. Without loss of generality, assume $Q=\mathbb{N}$. Define the following relations:
  • For each $a\in A$ and $c\in C$,
    $R_{ac}=\{(i,j)\mid i,j\in\mathbb{N},\;i\xrightarrow{a,c}j\};$
  • For each $a\in A$, define
    $G_{a}:\mathbb{N}\to\{\mathit{true},\mathit{false}\},$
    where $G_{a}(i)=\mathit{true}$ iff there exist $c\in C$ and $j\in\mathbb{N}$ such that $i\xrightarrow{a,c}j$;
  • Define
    $R_{Q_{0}}=\{(1,i)\mid i\in Q_{0}\}.$
Because $S$ is a computable controlled automaton, the relations $G_{a}$, $R_{ac}$, and $R_{Q_{0}}$ are computable. Hence, each can be represented by a nondeterministic Turing machine. By the universality result above, there exist corresponding activity networks
$K_{G_{a}},\quad K_{R_{ac}},\quad\text{and}\quad K_{Q_{0}},$
constructed solely from instantaneous activities, standard gates, and inhibitor gates, that simulate these relations.
We now construct a controlled activity network $K$ whose set of timed activities is $A$ and whose set of control actions is $C$. For each pair $(a,c)\in A\times C$, $K$ contains a controlled activity subnetwork of the form depicted in Figure 5. In this subnetwork, $K_{G_{a}}$ and $K_{R_{ac}}$ simulate the enabling predicate $G_{a}$ and the transition relation $R_{ac}$, respectively.
Let $P_{1a},P_{2a},P_{3ac}$, and $P_{M}$ be places with the following roles: when $P_{1a}$ is empty, all instantaneous activities of $K_{G_{a}}$ are disabled; when $P_{3ac}$ is empty, all instantaneous activities of $K_{R_{ac}}$ are disabled. Initially, $P_{1a}$ contains one token, while $P_{2a}$ and $P_{3ac}$ are empty.
Whenever the system state place $P_{S}$ acquires a marking $x$ such that $G_{a}(x)=\mathit{true}$, the subnetwork $K_{G_{a}}$ executes. After finitely many instantaneous firings, $P_{1a}$ loses its token, $P_{2a}$ gains a token, and the marking of $P_{S}$ remains unchanged. When $P_{3ac}$ subsequently gains a token, the subnetwork $K_{R_{ac}}$ executes. After finitely many instantaneous firings, $P_{3ac}$ loses its token, $P_{1a}$ regains a token, and the marking of $P_{S}$ changes from $x$ to $y$ such that $(x,y)\in R_{ac}$.
Initialization is handled by an additional controlled activity subnetwork, depicted in Figure 6. Here, $K_{Q_{0}}$ simulates the relation $R_{Q_{0}}$. Initially, a place $P_{0}$ contains one token and $P_{S}$ is empty. After finitely many instantaneous firings, $P_{0}$ loses its token and the marking of $P_{S}$ is set to some $x$ with $(1,x)\in R_{Q_{0}}$, after which all activities of $K_{Q_{0}}$ are disabled.
Assuming that $K$ contains no controlled activity subnetworks other than those described above, it follows that the controlled activity network $K$ realizes a controlled automaton that is isomorphic to $S$. This completes the proof. ∎